# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ Abstract
We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time effects in both the intercept and slope. To address the incidental parameter problem inherent in this class of models, we develop an analytical bias correction procedure to construct a bias-corrected likelihood. The resulting maximum likelihood estimators are automatically bias-corrected. Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-square distribution under the null hypothesis. Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
@@
incidental parameter problem; bias correction; fixed effects; maximum likelihood
@@ 1 Introduction
Panel data models with two-way fixed effects are widely used in empirical research to control for unobserved heterogeneity that may correlate with covariates. Individual effects capture time-invariant, individual-specific characteristics, while time effects account for shifts that are common across individuals but vary over time. Traditionally, these models address level heterogeneity by incorporating fixed effects only in the intercept, adjusting for baseline outcome differences across individuals and time periods. However, many empirical applications reveal substantial variation in how outcomes respond to covariates (Browning and Carro, 2007). When such response heterogeneity are present and correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
@@
In this paper, we introduce a general class of dynamic nonlinear heterogeneous parameter models (DN-HP) that incorporate individual and time fixed effects in both the intercept and slope. This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel data models. To address the incidental parameter problem inherent in DN-HP models, we propose a likelihood-based analytical bias correction approach. This method ensures the validity of the maximum likelihood estimator and likelihood-based inference for the structural parameters of DN-HP models. As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing maximum likelihood estimators from our bias-corrected likelihoods and the original likelihoods. The results reveal that the maximum likelihood estimators obtained from the bias-corrected likelihoods significantly reduce estimation bias while maintaining comparable mean square errors.
@@
Notes: Model. $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of $X_{it},$ and the true values $\rho_{0}=\beta_{0}=0.5$. ($\widehat{\beta},\widehat{\rho})$ are the uncorrected estimators and $(\widehat{\beta}_{L},\widehat{\rho}_{L})$ are the corrected estimators. $1000$ replications. For details on the data-generating process, please refer to Section 4.
1.1 Literature Review
The estimation and inference of fixed effects models in the presence of the incidental parameter problem have been extensively studied. Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, addressing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models. In recent decades, the availability of long panel datasets has motivated a large-$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow proportionally to infinity. Under this rectangular asymptotic, estimators remain consistent but typically exhibit a non-negligible bias of order $O(1/T)$ due to the presence of individual effects in the intercept. To enable valid inference, researchers have developed various bias correction methods for a broad range of models in a maximum likelihood framework. For instance, Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” bias correction methods to eliminate the bias from the maximum likelihood estimator. Two alternative approaches have also been explored: “score-based” and “likelihood-based” methods. Woutersen (2002) and Li et al. (2003) modify the profiled score to remove the bias, while Bester and Hansen (2009) and Arellano and Hahn (2016) identify and correct the bias directly for the likelihood. Most of these methods are analytical, estimating the bias using explicit formulas. Alternatively, numerical procedures are available, estimating the bias through methods such as the Jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023), and integrated likelihoods (Arellano and Bonhomme, 2009).
@@ 1.1 Literature Review
1.2 This Paper
We assume that the multi-dimensional fixed effects enter the log-likelihood additively. Identification restrictions are imposed on the fixed effects, whose dimension grows with the sample size. Our corrected likelihood is derived using two Taylor expansions: the expansion of the score function and the expansion of the original likelihood, both with respect to the fixed effects. Under appropriate dominance conditions, we explicitly derive the two additive bias terms in the original likelihoods of DN-HP models. These bias terms are expressed as quadratic forms involving the score and Hessian of the fixed effects. We estimate the derived bias terms under weak serial dependence (conditional strong mixing) and conditional cross-sectional independence.
@@ 1.2 This Paper
We rigorously show that the maximum likelihood estimator (MLE) based on the bias-corrected likelihood of the DN-HP model retains a limiting normal distribution with zero mean (automatically bias-corrected), consistent with the limiting distribution of the classical MLE in the absence of incidental parameters. Beyond the bias-corrected MLE, this likelihood-based bias correction approach also facilitates hypothesis testing using the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics. We demonstrate that these three test statistics, derived from the bias-corrected likelihood, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
@@
We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife. In the empirical application, we apply the DN-HP logit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households. Our analysis reveals that this impact varies significantly across individuals and over time. Neglecting such variation can lead to biased estimates and markedly different conclusions, highlighting the need to address both level and response heterogeneities, as well as the value of our framework in empirical research.
@@
The remainder of the paper is organized as follows. Section 2 describes our settings, the identification restriction, the incidental parameter problem under DN-HP, and gives an overview of our bias correction procedures. Section 3 provides the bias expansion of the likelihood and the formula for our corrected likelihood. The statistical properties of the corrected likelihood. We rigorously show that the estimator of the structural parameter, and the LR, LM, and Wald statistics all follow their standard asymptotic distribution. Section 4 presents simulation studies to demonstrate our method’s performance. Here we show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors. In addition, we also show that our procedure restores the test size of the LR test to its nominal level. We also compare our procedure with the Jackknife and find that our procedure outperforms the Jackknife. Section 5 applies our method to study single mothers’ labor force participation. Proofs, technical details, and additional results are provided in the appendix.
@@
Notations. In what follows, we denote $\iota_{n}$ to be an $n\times 1$ vector of ones and for $\mathbb{I}_{n}$ to be an $n\times n$ identity matrix. Let $\otimes$ be the kronecker product and $\mathds{1}\left\{\cdot\right\}$ the indicator function. In addition, for each square matrix $A_{i}$, $i=1,\ldots,n$, $\operatorname*{diag}\left\{A_{1},\ldots,A_{n}\right\}$ represent a block diagonal matrix, where each matrix $A_{i}$ is the $i$-th diagonal element.
2 Bias in Heterogeneous Parameter Model
2.1 DN-HP Model
Consider the responses $Y_{it},$ which follow the conditional density $f\left(Y_{it};\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}\right),$ given covariates $X_{it}$ (omitted for simplicity). Here, $i=1,\ldots,N$ represents individuals, and $t=1,\ldots,T$ represents time periods. The density is known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0},$ and the covariates $X_{it}$ may include lagged $Y_{it}$. Specifically, $\theta_{0}$ is the $K$-dimensional structural parameter of interest, where $K$ is fixed and known. The nuisance heterogeneous effects $\alpha_{i}^{0}=\left(\alpha_{1,i}^{0},\ldots,\alpha_{K,i}^{0}\right)^{\prime}$ and $\gamma_{t}^{0}=\left(\gamma_{1,t}^{0},\ldots,\gamma_{K,t}^{0}\right)^{\prime}$ may be correlated with $X_{it}$. Throughout the paper, we treat $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ as fixed effects (unknown parameters) by conditioning on them in the density.
Denote $\phi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime}\right)^{\prime}$ and define the quasi-likelihood function as
$l(\theta,\phi):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}\left(\theta,\phi\right),\qquad l_{i,t}(\theta,\phi):=\log f\left(Y_{it};\theta+\alpha_{i}+\gamma_{t}\right).$
Note that if the data are independent across both $i$ and $t,$ $l(\theta,\phi)$ represents the exact likelihood. However, when observations for each unit $i$ exhibit serial correlation, $l(\theta,\phi)$ is misspecified even though the likelihood for each observation $Y_{it}\mid X_{it},$ $l_{i,t}(\theta,\phi),$ is correctly specified. Our asymptotic theory assumes conditional cross-sectional independence and weak time-serial dependence (Assumption 1). For simplicity, we refer to $l(\theta,\phi)$ as the likelihood function hereafter. Here is an example that illustrates how our model captures greater heterogeneity than conventional two-way effects panel data models.

### v4

@@ Abstract
We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time fixed effects in both the intercept and slope. These models are subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at zero, and that test statistics do not follow their standard asymptotic distributions as in the absence of the fixed effects. To address the problem, we develop an analytical bias correction procedure to construct a bias-corrected likelihood. The resulting estimator follows an asymptotic normal distribution with mean zero. Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-squared distribution under the null hypothesis. Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
@@
incidental parameter problem; bias correction; fixed effects; panel data models
@@ 1 Introduction
Panel data is common in empirical research. Panel data models with two-way fixed effects are widely used to control for unobserved heterogeneity that may correlate with covariates. Individual effects capture time-invariant, individual-specific characteristics, while time effects account for shifts that are common across individuals but vary over time. Traditionally, these models address level heterogeneity by incorporating fixed effects only in the intercept, adjusting for baseline outcome differences across individuals and time periods. However, many empirical applications reveal substantial variation in how outcomes respond to covariates (Browning and Carro, 2007). When such response heterogeneity are present and correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
@@
In this paper, we introduce a general class of dynamic nonlinear heterogeneous parameter models (DN-HP) that incorporate individual and time fixed effects in both the intercept and slope. This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel data models. DN-HP models are generally subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at 00, and that the test statistics do not follow their standard asymptotic distributions (of $\chi^{2}$, for many test statistics). To address the incidental parameter problem, we propose an analytical bias correction approach. This method restores the standard asymptotic distributions for the point estimator and the test statistics.
@@
{justify}
Notes: Model. $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of $X_{it}$, and the true values $\rho_{0}=\beta_{0}=0.5$. $\widehat{\beta}$ and $\widehat{\rho}$ are the uncorrected estimators while $\widehat{\beta}_{L}$ and $\widehat{\rho}_{L}$ are the corrected estimators. $1000$ replications. The detailed data-generating process is given in Section 4.
As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing estimators from our bias correction procedure to the uncorrected ones. The results reveal that our bias correction procedure significantly reduces the estimation bias while maintaining comparable mean square errors.
Literature Review. The estimation and inference of fixed effects models in the presence of the incidental parameter problem have been extensively studied. Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, addressing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models. In recent decades, the availability of long panel datasets has motivated a large-$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow proportionally to infinity. Under this rectangular asymptotic, estimators remain consistent but typically exhibit a non-negligible bias of order $O(1/T)$ due to the presence of individual effects in the intercept. To enable valid inference, researchers have developed various bias correction methods for a broad range of models in a maximum likelihood framework. For instance, Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” bias correction methods to eliminate the bias from the maximum likelihood estimator. Two alternative approaches have also been explored: “score-based” and “likelihood-based” methods. Woutersen (2002) and Li et al. (2003) modify the profiled score to remove the bias, while Bester and Hansen (2009) and Arellano and Hahn (2016) identify and correct the bias directly for the likelihood. Most of these methods are analytical, estimating the bias using explicit formulas. Alternatively, numerical procedures are available, estimating the bias through methods such as the Jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023), and integrated likelihoods (Arellano and Bonhomme, 2009).
@@
This paper. We employ the maximum likelihood framework and focus on multi-dimensional additive two-way fixed effects, whose number grows with the sample size. We consider an arbitrary DN-HP model whose log-likelihood function is specified up to the unknown parameters. We construct a modified, or corrected, log-likelihood function by adding two bias correction terms to the original one. These two bias terms are analytically derived by combining i) the Taylor expansion of the original log-likelihood function, in terms of the fixed effects, and ii) an asymptotic expansion of the fixed effects themselves. We show that the corrected log-likelihood function does not suffer from the incidental parameter problem, under appropriate regularity conditions.
@@
One key benefit of our procedure is that we achieve bias corrections for the point estimators and the test statistics by modifying a single object, the log-likelihood function. We rigorously show that the estimators obtained by maximizing the corrected log-likelihood function retains a limiting normal distribution with zero mean, consistent with the asymptotic theory of the classical maximum likelihood estimators (MLE) in the absence of incidental parameters. Beyond the estimators, we also show that the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics, derived from the same corrected log-likelihood function, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
@@
We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife, which is a potential bias correction device [^1] It is worth mentioning that, to our knowledge, there currently has been no study on using the jackknife under the DN-HP models.. We show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors, and restores the test sizes of the LR test to its nominal level. We also find that our procedure outperforms the Jackknife. In the empirical application, we apply the DN-HP logit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households. Our analysis reveals that this impact varies significantly across individuals and over time. Neglecting such variation can lead to biased estimates and markedly different conclusions, highlighting the need to address both level and response heterogeneities, as well as the value of our framework in empirical research.
@@
The remainder of the paper is organized as follows. Section 2 explains our settings, the identification restriction, and the incidental parameter problem of DN-HP models. Section 3 describes our bias correction procedure and provides relevant statistical properties, including those we claim above. Section 4 presents simulation studies to demonstrate the performance of method. Section 5 applies our method to study single mothers’ labor force participation. Finally in Section 6, we leave some closing remarks. Proofs, technical details, and additional results are provided in the appendix.
@@
Notation. We denote $\mathbb{I}_{n}$ to be the $n\times n$ identity matrix and $\iota_{n}$ the $n\times 1$ vector of ones. $\otimes$ denotes the Kronecker product and $\mathds{1}\{\cdot\}$ is the indicator function. Next, for a sequence of square matrices $A_{i}$, $i=1,\ldots,n$, $\operatorname*{diag}\{A_{1},\ldots,A_{n}\}$ represents a block diagonal matrix where each $A_{i}$ is the $i$-th diagonal element. For a vector $v$ and a function $f(v)$, we write $\partial_{v}$ and $\partial_{vv^{\prime}}$ to represent the first and second partial derivatives of $f(v)$ with respect to (w.r.t.) $v$.
2 Bias in Heterogeneous Parameter Models
2.1 Models and Estimations
Let $\{(Y_{it},X_{it}^{\prime})^{\prime}:i=1,\ldots,N;t=1,\ldots,T\}$ be a panel data set of $N$ individuals and $T$ time periods, where, $i$ represents the individual and $t$ represents the time period. Here $Y_{it}$ is a scalar response variable, and $X_{it}$ is a $K\times 1$ vector of regressors, with $K$ known and fixed. For each individual $i$, the response $Y_{it}$ is generated sequentially over $t$, by
$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}\sim f(y|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$ (2.1)
where $\theta_{0}$ is a $K\times 1$ vector of (non-random) unknown common parameters of interests, $\phi^{0}:=(\alpha_{1}^{0\prime},\ldots,\alpha_{N}^{0\prime},\gamma_{1}^{0\prime},\ldots,\gamma_{T}^{0\prime})^{\prime}$ consists of the $K\times 1$ vectors of random variables $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, and $f(y|\cdot)$ is a density known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$. We restrict our attention to the case where $f\left(\cdot\right)$ depends on $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$ through the additive structure $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, the “heterogeneous parameter”. The vectors $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ capture the individual-specific and time-specific heterogeneities, respectively. Generally, $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ may be correlated with the regressor $X_{it}$. However, throughout the paper, we employ the “fixed-effect” framework to condition on the realizations of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ (denoted by the same symbols) in the density, and treat them as unknown nuisance parameters to be estimated together with $\theta_{0}$. Conditioning on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, we assume $(Y_{it},X_{it}^{\prime})^{\prime}$ independent across $i$ but serially dependent over $t$. In particular, we allow $X_{it}$ to contain both strictly exogenous components and predetermined components w.r.t. $Y_{it}$. For instance, $X_{it}$ may contain lagged values of $Y_{it}$. In this paper, we focus on the case where only heterogeneous parameters appear in (2.1). This is without loss of generality. As illustrated in the example below, our approach accommodates settings where both heterogeneous parameters and “homogeneous parameters” are present. [^2] Another example is the classic Neyman-Scott model, where $Y_{it}$ follows a normal distribution with a homogeneous variance but heterogeneous means. See Example 3 below.